Consider the experiment of rolling a pair of dice. Let X be the random variable representing the absolute value of the difference of the numbers that come up






A) Construct the probability mass function (Tabular, Formula and Graphical Form).

1
Expert's answer
2022-03-16T16:04:34-0400

solution

Total number of cases when two dice roles =6×6=36


The absolute values of the difference of the numbers that come up are as follows

"\begin{matrix} &1 & 2 & 3 & 4 & 5 & 6 \\ 1 & 0 & 1 & 2& 3 & 4& 5 \\ 2& 1 & 0& 1& 2 & 3& 4\\ 3 & 2 & 1 & 0& 1 & 2& 3\\ 4 & 3 & 2& 1 & 0 & 1 & 2\\ 5 & 4 & 3 & 2& 1 & 0& 1\\ 6 & 5 & 4 & 3 & 2 & 1 & 0 \end{matrix}"


The probability of each value is as shown

(a) Tabular